### 17 Jan 2024 v3

# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

Abstract
[…]
1 Introduction
Multivariate time series (TS) forecasting entails predicting future values for multiple interrelated time series based on their historical data. This field has advanced significantly, applying statistical and machine learning (ML) methods Hyndman and Athanasopoulos (2021) across domains like weather, traffic, retail, and energy. In general, each time series represents a variable or channel111“Channel” refers to the individual time series in multivariate data (i.e., a multivariate TS is a multi-channel signal).. In certain applications, non-forecasting variables, categorized as controllable and uncontrollable external factors, impact the variables to forecast. We term these non-forecasting variables as exogenous, and the variables requiring forecast as target variables.
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al. (2017) approaches, exemplified by models like PatchTST Nie et al. (2023), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDFormer Zhou et al. (2022). These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer Tolstikhin et al. (2021), such as TSMixer Ekambaram et al. (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training. There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al. (2023) and TF-C Zhang et al. (2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al. (2023) and GPT4TS Zhou et al. (2023) yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
However, these LLM based TS approaches do not explicitly handle channel correlations and exogenous support in the context of multivariate forecasting. Moreover, these large models, with billions of parameters, demand significant computational resources and runtime. Hence, in this paper, we focus on building pre-trained models from scratch solely using TS data. Unlike language, which has abundant public pre-training data in terabytes, time-series data is relatively scarce, very diverse and publicly limited. Its scarcity leads to overfitting when pre-training “large” models solely on time-series data. This prompts a question: Can smaller models pre-trained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy? Surprisingly, the answer is yes! Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al. (2021)). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pre-training on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pre-trained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
Below, we outline the paper’s key contributions:
  • Amidst the prevalence of large pre-trained models demanding significant compute and training time (in weeks), our work is the first to showcase the efficacy of building Fast and Tiny Pre-trained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just few hours (4-8 hours, 6 A100 GPUs). TTM successfully demonstrates transfer learning to diverse, unseen target datasets for zero/few-shot forecasting, addressing the data scarcity issues prevalent in time series.
  • Pre-training on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models. Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pre-training.
  • TTM employs a multi-level modeling strategy to explicitly model channel-correlations, and incorporates exogenous signals – a crucial capability lacking in LLMs-based TS approaches.
  • With extensive evaluation on 11 datasets, TTM shows significant accuracy gains over popular benchmarks (12-38% in few/zero-shot forecasting). It also drastically reduces the compute needs as compared to LLM-TS methods, with a 14X cut in learnable parameters, 106X less total parameters, and substantial reductions in finetuning (65X), inference time (54X), and memory usage (27X).
  • The zero-shot results of TTM often surpass the few-shot results of many SOTA approaches, highlighting the effectiveness of our approach.
Figure 1: Overview of Multilevel Tiny Time Mixers (TTM): (a) Refer to Section 2 and 3, (b) Refer to Section 3.1, (c) Refer to Section 3.2
2 TTM Components
Let $\bm{X}\in\mathbb{R}^{c\times sl}$ be a multivariate time series of length $sl$ and $c$ number of channels. The forecasting task can be formally defined as predicting the future values $\bm{Y}\in\mathbb{R}^{c^{\prime}\times fl}$ given the history $\bm{X}$. Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes number of forecast channels, where $c^{\prime}\leq c$. The predictions from the model are denoted by $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g. weather), and (ii) control variables whose future values during the forecast horizon can be manipulated to govern the behavior of the target variables. (e.g. discount in sales forecasting, operator controls in industrial applications). In TTM, uncontrolled and control variables are treated similarly, as both are considered available during forecasting.
2.1 Multi-level Modeling
TTM follows a multi-level architecture consisting of four key components (see Figure 1(a)): (1) The TTM Backbone is assembled using building blocks derived from the efficient TSMixer architecture Ekambaram et al. (2023). TSMixer is based on simple MLP blocks that enable mixing of features within patches, across patches and channels, surpassing existing transformer-based TS approaches with minimal computational requirements. Since TSMixer is not targeted to handle multi-resolution data, we introduce various novel enhancements to it as explained later. (2) TTM Decoderfollows the same backbone architecture but is considerably smaller in size, approximately 10-20% of the size of the backbone, (3) Forecast Headconsists of a linear head designed to produce the forecast output, and (4) Optional Exogenous Mixer serves to fuse exogenous data into the model’s forecasting process. This multi-level model refactoring is required to dynamically change the working behavior of various components based on the workflow type, as explained in Section 3. In addition to the above primary components, we also have a preprocessing component as explained next.
2.2 Pre-processing
As shown in Figure 1(a) with colorless blocks, the historical time series $\bm{X}$ is first normalized per instance to have zero mean and unit standard deviation for each channel dimension, to tackle any possible distribution shifts Nie et al. (2023); Ekambaram et al. (2023). This process is reversed at the end before computing the loss. The normalized data $\overline{\bm{X}}$ is subsequently patched $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ into $n$ non-overlapping windows, each of length $pl$ and then, passed to the TTM backbone. Patching, as introduced in Nie et al. (2023), has proven to be highly valuable for forecasting. Its effectiveness lies in preserving local semantic information, accommodating longer history, and reducing computation.
3 TTM Workflows
TTM works in 2 stages: pre-train and fine-tune (Figure 1(a)).
3.1 Pre-training Workflow
In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al. (2021). Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective. TTM is first pre-trained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pre-training datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pre-training datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pre-training. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pre-train the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
3.1.1 Multi-Resolution Pre-training via TTM Backbone
The majority of the pre-training happens in the TTM backbone. The primary challenge with the proposed pre-training technique is that the pre-training data is diverse and has multiple resolutions. There are two main options for pre-training: conducting separate pre-training for each resolution type or pre-training using all resolution data collectively. While it’s common to train a model per resolution type to overcome challenges in learning diverse seasonal patterns, this leads to diminished training data for each resolution due to limited data availability. Consequently, this motivated the exploration of pre-training a single model using datasets from all resolutions. To achieve this, we propose the following 3 enhancements.
Adaptive Patching: The TTM backbone is crafted with an adaptive patching architecture where different layers of the backbone operate at varying patch lengths and numbers of patches. Since each dataset in the pre-training corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced. As shown in Figure 1(b), the patched data $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ is passed through a embedding layer to project it to the patch hidden dimension, $\bm{X}_{h}\in\mathbb{R}^{c\times n\times hf}$. Optionally, if the resolution prefix tuning module is activated (as explained later), the resolution prefix is concatenated with $\bm{X}_{h}$. For notational simplicity, we denote the concatenated tensor with $\bm{X}_{h}$ as well.
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block. Each TTM block is further comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pre-train with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers Liu et al. (2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$. Please note that adaptive patching is enabled only in the backbone and not in the decoder, which is designed to be very lightweight.
Data Augmentation via Downsampling: A significant challenge in TS pre-training datasets is the scarcity of public datasets at specific resolutions. To overcome this, we employ a downsampling technique for high-resolution datasets, generating multiple datasets at lower resolutions. For example, from a one-second resolution dataset, we derive datasets at minute and hour resolutions. Note that, the original high-resolution dataset remains within the pool of pre-training datasets. This methodology significantly augments the number of datasets for each resolution which greatly improves the model performance (Section 4.5).
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In these scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
3.2 Fine-tuning Workflow
In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pre-training datasets. We have three options here: (a) In Zero-shot forecasting, we directly use the pre-trained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pre-trained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we fine-tune the pre-trained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
The backbone is completely frozen during fine-tuning, and still operates in a channel-independent univariate fashion. However, the TTM decoder can be fine-tuned via channel-mixing (for multivariate) or a channel-independent (for univariate) way based on the nature of the target data. If pure multivariate modeling is needed, then the channel-mixer block in all the TSMixer components (see Figure 1(b)) in the decoder gets enabled to explicitly capture the channel correlation between the channels. The forecast head and reverse normalization perform similar operations as in the pre-training stage. The fine-tuning also optimizes the forecasting objective with MSE loss. This thoughtful multi-level design choice ensures our backbone excels in channel-independent pre-training, enabling effective temporal correlation modeling across diverse datasets. Simultaneously, the decoder handles target-data-specific tasks like channel-correlation modeling and fine-tuning. In addition, if the target data has exogenous variables, then an exogenous mixer block is applied to the actual forecasts as explained next.
Exogenous Mixer Block: As described in Section 2, the future values of the exogenous channels are known in advance. Let the forecast from the forecast head be $\hat{\bm{Y}}\in\mathbb{R}^{c\times fl}$. Let the channels $\bm{x}_{0},\cdots,\bm{x}_{c^{\prime}}$ denote the target variables and $\bm{x}_{c^{\prime}+1},\cdots,\bm{x}_{c}$ denote all exogenous variables with their future values known. First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{c^{\prime}},\bm{y}_{c^{\prime}+1},\cdots,\bm{y}_{c}\right]\in\mathbb{R}^{fl\times c}$. Next, to learn inter-channel lagged correlations, we patch $\hat{\bm{Y}}_{e}$ into a series of overlapped windows (i.e., patching with stride$=1$) to create a new tensor: $\hat{\bm{Y}}_{e,p}\in\mathbb{R}^{fl\times\Delta\times c}$, where $\Delta=2\cdot l+1$ with $l$ being the context length to incorporate on either side of a time point [^4] This needs padding $\hat{\bm{Y}}_{e}$ with zeros of length $l$ on both sides.. Subsequently, we pass $\hat{\bm{Y}}_{e,p}$ through a vanilla TSMixer block with channel mixing enabled. Thus, the lagged dependency of the forecasts for the target channels on the exogenous channels is seamlessly learned. Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. Figure 1(c) depicts this procedure.
Zero-shot Few-shot (5%) $fl$ TTM TTM GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 96 0.365 0.366 0.543 0.557 0.554 0.892 0.547 0.593 0.681 192 0.393 0.391 0.748 0.711 0.673 0.940 0.720 0.652 0.725 336 0.415 0.421 0.754 0.816 0.678 0.945 0.984 0.731 0.761 720 0.538 - - - - - - - - ETTH2 96 0.285 0.282 0.376 0.401 0.348 0.409 0.442 0.390 0.428 192 0.341 0.338 0.418 0.452 0.419 0.483 0.617 0.457 0.496 336 0.383 0.383 0.408 0.464 0.389 0.499 1.424 0.477 0.486 720 0.441 - - - - - - - - ETTM1 96 0.413 0.359 0.386 0.399 0.361 0.606 0.332 0.628 0.726 192 0.476 0.402 0.440 0.441 0.411 0.681 0.358 0.666 0.750 336 0.553 0.424 0.485 0.499 0.467 0.786 0.402 0.807 0.851 720 0.737 0.575 0.577 0.767 0.677 0.796 0.511 0.822 0.857 ETTM2 96 0.187 0.174 0.199 0.206 0.200 0.220 0.236 0.229 0.232 192 0.261 0.240 0.256 0.264 0.265 0.311 0.306 0.394 0.291 336 0.323 0.299 0.318 0.334 0.314 0.338 0.380 0.378 0.478 720 0.436 0.407 0.46 0.454 0.410 0.509 0.674 0.523 0.553 Weather 96 0.154 0.152 0.175 0.171 0.188 0.207 0.184 0.229 0.227 192 0.203 0.198 0.227 0.230 0.234 0.272 0.228 0.265 0.278 336 0.256 0.250 0.286 0.294 0.287 0.313 0.279 0.353 0.351 720 0.329 0.326 0.366 0.384 0.365 0.400 0.364 0.391 0.387 Electricity 96 0.169 0.142 0.143 0.145 0.147 0.315 0.150 0.235 0.297 192 0.196 0.162 0.159 0.163 0.172 0.318 0.163 0.247 0.308 336 0.209 0.184 0.179 0.175 0.190 0.340 0.175 0.267 0.354 720 0.264 0.228 0.233 0.219 0.280 0.635 0.219 0.318 0.426 Traffic 96 0.518 0.401 0.419 0.404 0.408 0.854 0.427 0.670 0.795 192 0.548 0.425 0.434 0.412 0.421 0.894 0.447 0.653 0.837 336 0.55 0.437 0.449 0.439 0.477 0.853 0.478 0.707 0.867 720 0.605 - - - - - - - - TTM Zero-shot % IMP 1% 3% 2% 31% 6% 25% 32% TTM 5% Few-shot % IMP 12% 14% 12% 38% 17% 32% 38% Table 1: Zero-shot and Few-shot 5% performance (MSE) of TTM and all SOTA models on D1 datasets for varying forecast lengths ($fl$). Bold and underscore denote the best and the second best. For longer $fl$ (720), some datasets fail to generate fine-tuning samples with just 5% of the data, denoted as ‘-’. The last two rows show relative improvement (IMP) of TTM zero-shot and few-shot w.r.t. SOTA few-shot results.
Metric Data GPT4TS TTM (Ours) IMP (X) FT NPARAMS (M) Weather 3.2 0.29 (14X) Electricity 3.2 0.29 Traffic 5.5 0.29 TL NPARAMS (M) Weather 84.3 0.8 (106X) Electricity 84.3 0.8 Traffic 86.6 0.8 EPOCH TIME (s) Weather 37 1 (65X) Electricity 241 4 Traffic 782 8 MAX MEMORY (GB) Weather 21.9 1.06 (27X) Electricity 39.7 2.02 Traffic 53.7 1.34 TEST TIME (s) Weather 52.53 1.66 (54X) Electricity 353.4 8.07 Traffic 1163.9 13.5 Table 2: Computational Improvement in few-shot 10% for $fl$=96 using 1 A100 GPU. nX denotes n times average improvement across datasets (IMP). 10% 25% 50% 75% 100% IMP TTM 0.422 0.421 0.413 0.402 0.398 - SimMTM 0.591 0.535 0.491 0.466 0.415 17% Ti-MAE 0.660 0.594 0.55 0.475 0.466 24% TST 0.783 0.641 0.525 0.516 0.469 28% LaST 0.645 0.610 0.540 0.479 0.443 23% TF-C 0.799 0.736 0.731 0.697 0.635 43% CoST 0.784 0.624 0.540 0.494 0.428 26% TS2Vec 0.655 0.632 0.599 0.577 0.517 31%
Figure 2: Cross transfer learning MSE improvement (IMP) for self-supervised pre-training methods in various few-shot settings (10%,25%,50%,75%,100%).
4 Experiments and Results
4.1 Experimental Setting
Datasets: Pre-training employs a subset of the Monash data hub Godahewa et al. (2021) of size $\sim 244$M samples. We specifically exclude datasets (like yearly, monthly) as they do not possess sufficient length for the long-term forecasting task. Moreover, we remove all the datasets that we utilize for evaluation (i.e., weather, electricity, and traffic). For zero/few-shot evaluation we consider seven public datasets (D1): ETTH1, ETTH2, ETTM1, ETTM2, Weather, Electricity & Traffic as popularly used in most prior SOTA works Zhou et al. (2021); Nie et al. (2023). Since these datasets do not contain any exogenous variables nor exhibit cross-channel correlation benefits, we incorporate four other datasets (D2) for separately validating the efficacy of the decoder channel mixing and exogenous mixer module: bike sharing (BS) Fanaee-T (2013), carbon capture plant (CC) Jablonka et al. (2023), and 2 more datasets from Biz-IT observability domain ITB (2023): Application (APP) and Service (SER). Refer Appendix for full data details.
SOTA Bechmarks: We benchmark TTM with the latest public SOTA forecasting models categorized as follows: [label=()] 1. LLM-based TS pre-trained models:GPT4TS Zhou et al. (2023), LLMTime Gruver et al. (2023) 2. Self-supervised pre-trained models: SimMTM Dong et al. (2023),Ti-MAE Li et al. (2023), TST Zerveas et al. (2021), LaST Wang et al. (2022), TF-C Zhang et al. (2022), CoST Woo et al. (2022) and Ts2Vec Yue et al. (2022) 3. TS transformer models:PatchTST Nie et al. (2023), FEDFormer Zhou et al. (2022), Autoformer Wu et al. (2021) 4. Other SOTA models:TSMixer Ekambaram et al. (2023), DLinear Zeng et al. (2022) and TimesNet Wu et al. (2022)
Zero-shot Few-shot (10%) Data TTM (Ours) TTM (Ours) GPT4TS PatchTST TSMixer TimesNet Dlinear FEDFormer Autoformer ETTH1 0.428 0.422 0.590 0.633 0.578 0.869 0.691 0.639 0.702 ETTH2 0.362 0.360 0.397 0.415 0.370 0.479 0.605 0.466 0.488 ETTM1 0.545 0.458 0.464 0.501 0.441 0.677 0.411 0.722 0.802 ETTM2 0.302 0.278 0.293 0.296 0.282 0.320 0.316 0.463 1.342 Electricity 0.210 0.180 0.176 0.176 0.180 0.323 0.180 0.346 0.431 Traffic 0.555 0.428 0.440 0.430 0.429 0.951 0.446 0.663 0.749 Weather 0.236 0.227 0.238 0.242 0.236 0.279 0.241 0.284 0.300 TTM Zero-shot improvements -4% -2% -7% 27% 2% 27% 39% TTM Few-shot (10%) improvements 7% 9% 4% 34% 13% 34% 45% Table 3: Zero-shot and Few-shot 10% performance (MSE) of TTM over all SOTA models on D1 datasets, averaged across 4 forecast lengths. The numbers with bold and underscore denote the best and the second best results respectively. The full table is in Appendix.
Data $fl$ Zero-shot LLMTime TTM Traffic 96 0.340 0.195 192 0.526 0.367 Weather 96 0.107 0.031 192 0.062 0.067 Electricity 96 0.609 0.238 192 0.960 0.488 ETTM2 96 0.167 0.164 192 0.198 0.228 TTM IMP (%) 29% Table 4: TTM vs LLMTime MSE Improvement (IMP) in zero-shot setting.
TTM Model Details: For our experiments, we build 5 pre-trained models for the following $sl$,$fl$ configuration: (512,96), (512,192), (512, 336), (512,720) and (96,24). Based on the $sl$ and $fl$ requirement of the target dataset, a suitable pre-trained model is selected to initialize the weights accordingly for fine-tuning. Since our pre-trained models are very small, they can be trained quickly in a few hours (4-8 hrs), as opposed to several days/weeks in standard approaches. Hence, pre-training multiple TTM models is no longer a practical constraint. Pre-training is performed in a distributed fashion with 50 CPUs and 6 A100 GPUs with 40 GB GPU memory while fine-tuning uses only 1 GPU. Other model configurations are as follows: patch length $pl$ = 64 (for $sl$ 512) and 8 (for $sl$ 96), stride $s$ = $pl$, number of levels $L$ = 6, number of TTM blocks per level $M$ = 2, number of decoder layers = 2, batch size $b$ = 3K, number of epochs $ep$ = 20, and dropout $do$ = 0.2. TSMixer specific hyperparameters include feature scaler $fs$ = 3, hidden feature size $hf$ = $fs*pl$, expansion feature size $ef=hf*2$. Resolution prefix tuning is disabled by default and enabled only for shorter context lengths (as explained in Table. 9). Decoder channel-mixing and exogenous mixer blocks are disabled during pre-training and enabled for D2 datasets during fine-tuning. All other model parameters remain the same across pre-training and fine-tuning except for a few parameters like dropout and batch size that can be adjusted based on the target data. Hyperparameters are selected based on the validation performance, and the test results are reported. For full implementation details, refer to the Appendix. MSE is used as the default evaluation metric.
BS CC APP SER % IMP
TTM-CM 0.582 0.250 0.042 0.114
TTM-Zero-shot 0.992 0.263 0.183 0.238 44%
TTM 0.635 0.261 0.073 0.143 18%
PatchTST 0.735 0.267 0.060 0.119 15%
TSMixer-CC 0.651 0.284 0.053 0.136 15%
TSMixer-CM 0.716 0.303 0.069 0.118 20%
TSMixer 0.664 0.267 0.066 0.134 17%
GPT4TS 0.645 0.254 0.075 0.135 18%
Table 5: Effect of Decoder Mixing and Exog Fusion. MSE Results are reported using $(sl,fl)$ with values of $(512,96)$ for BS dataset and $(96,24)$ for other D2 datasets.
4.2 TTM’s Zero/Few-shot Performance
Table 1 compares the performance of TTM model in zero-shot and few-shot (5%) settings across multiple $fl$s. Baseline results are reported from  Zhou et al. (2023) as we use the same few-shot data filtering strategy as followed in that paper. Note that for zero-shot performance, the pre-trained TTM model is directly evaluated on the test set. The 5% few-shot TTM outperforms the SOTAs in most of the cases with significant accuracy gains (12-38%). An even more impressive observation is that the TTM in zero-shot setting is also able to outperform most of the SOTAs which are trained on 5% of the target data. This observation establishes the generalization ability of the pre-trained TTM model on the target datasets. Likewise, Table 3 shows the 10% few-shot performance of the TTM, where we outperform all the existing SOTAs by 4-45% accuracy gains. In addition, TTM zero-shot also beats many SOTAs (not all) with 10% training highlighting the effectiveness of our approach.
Additionally, we conduct a comparison between TTM and the LLMTime model which is explicitly designed for zero-shot setting. Since these models are based on LLaMA and are massive in size, the authors used only the last window of the standard test-set for faster evaluation, as opposed to using all the windows in the full test-set . Hence, we compare them separately in Table 4 based on the same datasets, $fl$, and test-set as reported in  Gruver et al. (2023). In this evaluation, we outperform LLMTime by a substantial margin of 29%. In addition, there are alternative pre-training approaches like masked modeling and contrastive learning techniques that may not offer universal transfer learning capabilities across all datasets (like TTM), but they excel in enabling cross-transfer learning between two datasets when carefully selected. Table 2 illustrates the cross-transfer learning from ETTH2 to ETTH1 for these models in various few-shot settings (as reported in  Dong et al. (2023)). Notably, TTM, with no specific cross-data selection, outperforms all popular SOTAs, including the latest SimMTM, by 17-43%. Thus, TTM significantly beats all the existing SOTAs including the recent popular transfer learning benchmarks based on LLMs. Notably, we achieve these accuracy gains with significantly reduced compute requirements, as elaborated next.
Data Zero-shot Few-shot (5%)) Few-shot (10%))
5 Epochs 50 Epochs 5 Epochs 50 Epochs
RI PT RI PT RI PT RI PT RI PT
ETTH1 0.735 0.428 0.687 0.389 0.629 0.393 0.64 0.419 0.552 0.422
ETTH2 0.400 0.362 0.378 0.334 0.379 0.334 0.371 0.357 0.370 0.360
ETTM1 0.722 0.545 0.474 0.483 0.448 0.440 0.421 0.458 0.403 0.458
ETTM2 0.353 0.302 0.296 0.274 0.290 0.280 0.279 0.271 0.274 0.278
Electricity 0.890 0.210 0.356 0.188 0.199 0.179 0.220 0.186 0.178 0.180
Traffic 1.453 0.555 0.593 0.454 0.441 0.421 0.486 0.446 0.430 0.428
Weather 0.321 0.236 0.267 0.233 0.264 0.231 0.240 0.229 0.238 0.227
IMP 36% 21% 12% 9% 2%
Table 6: Improvement (IMP) of Pre-Training (PT) over Randomly Initialized (RI) weights. MSE averaged across all $fl$s.
Data Zero-shot Few-shot (10%) Vanilla Adaptive Patching Vanilla Adaptive Patching ETTH1 0.501 0.428 0.478 0.422 ETTH2 0.375 0.362 0.372 0.360 ETTM1 0.577 0.545 0.463 0.458 ETTM2 0.310 0.302 0.272 0.278 Electricity 0.215 0.210 0.178 0.180 Traffic 0.556 0.555 0.431 0.428 Weather 0.239 0.236 0.230 0.227 IMP (%) 4% 2% Table 7: Effect of adaptive Patching, MSE averaged across 4 values of $fl$
Data Zero-shot Original Data With Downsample ETTH1 0.603 0.379 ETTH2 0.351 0.313 ETTM1 0.852 0.444 ETTM2 0.261 0.224 Electricity 0.424 0.182 Traffic 0.898 0.533 Weather 0.186 0.178 IMP (%) 30% Table 8: Downsampling effect, MSE averaged across $fl$s: 96, 192
Data $HL$ = 96 $HL$ = 512 NP P NP P ETTH1 0.373 0.358 0.365 0.360 ETTH2 0.180 0.179 0.285 0.280 ETTM1 0.559 0.387 0.413 0.384 ETTM2 0.127 0.108 0.187 0.194 Weather 0.103 0.103 0.154 0.160 Electricity 0.208 0.201 0.169 0.175 Traffic 0.754 0.740 0.518 0.518 IMP (%) 8% 0% Table 9: Resolution Prefix Tuning:With (P) vs Without(NP). MSE reported
4.3 Computational Benefits of TTM
Table  2 compares the computational benefits of TTM over GPT4TS, the most popular LLM-TS model and the current best SOTA in few-shot setting. For a fair comparison, we run both models with the best-reported parameters in a single A100 GPU with the same setup (Details in Appendix). Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution. In specific, TTM achieves 14X cut in Fine-tune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters. This reduction in model footprint further leads to a significant reduction in fine-tuning EPOCH TIME by 65X, MAX MEMORY usage by 27X, and total inference time on the entire test data (TEST TIME) by 54X. The computational benefits also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2. In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution.
4.4 TTM’s effectiveness in cross-channel and exogenous modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 5 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the performance of TTM-CM with the pre-trained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Other SOTAs are not reported here, considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al. (2023) as they are the latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
4.5 Ablation Studies
Effect of Pre-training: Table 6 illustrates the advantages of the proposed pre-training (PT) approach in comparison to randomly initialized (RI) model weights. In the case of zero-shot, the pre-trained TTM (PT) exhibits 36% improvement over RI. This outcome is expected, as random weights are directly employed for forecasting with RI. For the 5% few-shot scenario, PT achieves noteworthy improvements of 21% and 12% over RI when fine-tuned for 5 and 50 epochs, respectively. This underscores the utility of pre-trained weights in facilitating quick learning with a limited amount of target data. Even in the case of 10% few-shot, PT continues to demonstrate an improved performance of 9% for 5 epochs and 2% for 50 epochs. Thus, the performance impact of leveraging pre-trained weights greatly increases as the size of the training data and the available time-to-fine-tune reduces.
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone. Table 7 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. Performance improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
Effect of Augmentation via Downsampling: Table 8 compares the zero-shot test performances between the TTM model trained on the original Monash data and the TTM trained on the augmented Monash data after downsampling as described in Section 3.1. There is a significant improvement of 30% for the model pre-trained on the augmented datasets as it adds more data coverage across various resolutions. This observation highlights the strength of the proposed TTM model to learn from diverse datasets together, and increasing the number of datasets for each resolution helps.
Effect of Resolution Prefix Tuning: Since resolution prefix tuning explicitly adds the resolution type as an extra embedding, it greatly helps when the input context length $sl$ is short where it is challenging for the model to derive the resolution type automatically. As observed in Table 9, we observe 8% improvement in zero-shot for shorter context length ($sl$ = 96) and no improvement when the context length is longer ($sl$ =512). Hence this component is optional and can be enabled when working with shorter input context length.
5 Conclusions and Future Work
Considering the diversity and limited availability of public TS pre-training data, pre-training large models for effective transfer learning poses several challenges in time series. Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pre-training on limited diverse multi-resolution datasets. TTM achieves state-of-the-art results in zero/few-shot forecasting, offering significant computational efficiency and supporting cross-channel and exogenous variables — critical features lacking in existing popular methods. Going forward, we plan to extend our approach to many other downstream tasks beyond forecasting for a purely foundational approach in time series.